OpenStudy (anonymous):

Suppose the general term of a sequence is T(n)=n^(2)-10n+18. (a) Which two terms of the sequence are equal to 2? Ans: 2nd and 8th (GOT IT) >>>>(b) Is 200 a term in the sequence?

OpenStudy (anonymous):

@Peter14

OpenStudy (anonymous):

so you're looking for 200=n^2-10n+18 solve it like a quadratic equation

OpenStudy (anonymous):

n=19.38749457 OR -9.38749457

OpenStudy (anonymous):

for a sequence you can only use integral values of n

OpenStudy (anonymous):

so does the sequence have a term=200?

OpenStudy (anonymous):

ummm... substitute n=19/20??

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

errr....

OpenStudy (anonymous):

is it the factor=the term 200?

OpenStudy (anonymous):

when I learned about sequences the teacher said that you can only use integral values of n

OpenStudy (anonymous):

i know

OpenStudy (anonymous):

since the two values that would make x^2-10x+18=200 are not integral, 200 is not a term in the sequence. does that make any sense?

OpenStudy (anonymous):

ah, so if n is integer, then 200 is the answer?

OpenStudy (anonymous):

if you can get 200 by putting an integer in for n, 200 would be a term in the sequence.

OpenStudy (anonymous):

oh, i understood now:) thanks

OpenStudy (anonymous):

anything else?

OpenStudy (anonymous):

yup, about trigonometry